# Gauges and Accelerated Optimization over Smooth and/or Strongly Convex Sets

### v3

@@ A.3.1 Subgradient Method Convergence Theory
provided some minimizer $y^{*}$ of (20) has $\|y_{0}-y^{*}\|\leq D$. If additionally, each $\frac{1}{2}h_{i}^{2}$ is $\mu$-strongly convex, then selecting $s_{k}=\frac{2}{\mu(k+2)+\frac{M^{4}}{\mu(k+1)}}$ has
@@
where $\mathcal{D}(t)=Mp_{*}t+M^{2}t^{2}/2$. Then Theorem 1.2 of [51] ensures the subgradient method with normalized stepsizes $D/\sqrt{T+1}$ as we consider convergence at a rate of
@@
$\|h_{i}(y)g_{i}\|^{2}\leq M^{2}h^{2}(y)=M^{2}p_{*}^{2}+2M^{2}\left(\frac{1}{2}h^{2}_{i}(y)-\frac{1}{2}p_{*}^{2}\right).$ (45)
Since every subgradient of $h(y)$ is a convex combination of subgradients of the individual $h_{i}$, the same bound holds for each subgradient of $f$. Then in the strongly convex case, Theorem 1.7 of [51] ensures convergence for the average of the iterates at a rate of
@@ A.3.2 Accelerated Generalized Gradient Method Convergence Theory
We consider accelerated methods iteratively applying the “generalized gradient step” (26) following the development of Nesterov [17] (this can also be viewed as a prox linear step [32] on the composition of the maximum function $\max\{t_{1},\dots t_{m}\}$ with the smooth mapping $y\mapsto\frac{1}{2}(h^{2}_{1}(y),\dots,h^{2}_{m}(y))$). Computing this step can be formulated as a quadratic program of dimension $m$. When $m=1$, this is exactly a (sub)gradient step on $\frac{1}{2}h^{2}$ with stepsize $\alpha$. For $m=2$, this can be computed in closed form as
[…]
@@ A.3.3 Level Projection Method Convergence Theory
Finally, we consider the level-projection method (25). When $m=1$, (24) with $\bar{h}=p_{*}$ is exactly a (sub)gradient step with the Polyak stepsize rule. When $m=2$, this level projection step can still be computed in closed form:
@@
Let $\mathcal{\bar{Y}}=\{y\mid\frac{1}{2}h^{2}(y)\leq\frac{1}{2}\bar{h}^{2}\}\neq\emptyset$ denote the target level set, and $\mathcal{Y}_{k}=\{y\mid m_{k}(y)\leq\frac{1}{2}\bar{h}^{2}\}\supseteq\mathcal{\bar{Y}}$ denote our model’s target level set (the containment follows from $m_{k}$ lower bounding $\frac{1}{2}h^{2}$). Each orthogonal projection step ensures $x_{x}-x_{k+1}$ is normal to $\mathcal{Y}_{k}$. Consequently every $\bar{y}\in\mathcal{\bar{Y}}\subseteq\mathcal{Y}_{k}$ must have $(y_{k}-y_{k+1})^{T}(y_{k+1}-\bar{y})\leq 0$. Then
$\|y_{k+1}-\bar{y}\|^{2}$ $=\|y_{k}-\bar{y}\|^{2}+2(y_{k}-y_{k+1})^{T}(y_{k+1}-\bar{y})-\|y_{k+1}-y_{k}\|^{2}$
@@
Note that this recurrence implies $D_{k}$ is monotonically decreasing. Let $K$ denote the first index with $D_{k}\leq\bar{h}^{2}/\mu$. All $k<K$ must have a geometric decrease of $D_{k+1}\leq(1-\mu/8M^{2})D_{k}$. Thus $K\leq(8M^{2}/\mu)\log(\mu D_{0}/\bar{h}^{2})$. For all $k\geq K$, the recurrence ensures $D_{k+1}\leq D_{k}-\mu^{2}D_{k}^{2}/8M^{2}\bar{h}^{2}$, which implies $D_{k}\leq 8M^{2}\bar{h}^{2}/\mu^{2}(k-K+1)$. By assumption, we have $T/2>2K$ and so $T/2-K+1\geq T/4+1$. Hence recurrence ensures

### v4

@@ A.3.1 Subgradient Method Convergence Theory
provided some minimizer $y^{*}$ of (24) has $\|y_{0}-y^{*}\|\leq D$. If additionally, each $\frac{1}{2}h_{i}^{2}$ is $\mu$-strongly convex, then selecting $s_{k}=\frac{2}{\mu(k+2)+\frac{4M^{4}}{\mu(k+1)}}$ has
@@
where $\mathcal{D}(t)=Mp_{*}t+M^{2}t^{2}/2$. Then Theorem 1.2 of [53] ensures that the subgradient method with normalized stepsizes $D/\sqrt{T+1}$ achieves convergence at a rate of
@@
$\|h_{i}(y)g_{i}\|^{2}\leq M^{2}h_{i}^{2}(y)=M^{2}p_{*}^{2}+2M^{2}\left(\frac{1}{2}h^{2}_{i}(y)-\frac{1}{2}p_{*}^{2}\right).$ (53)
Since every subgradient of $h(y)$ is a convex combination of subgradients of the individual $h_{i}$, the same bound holds for each subgradient of $h$. Then in the strongly convex case, Theorem 1.7 of [53]  [^7] The proof of Theorem 1.7 of [53] gives a stronger statement which has been ultilized here. ensures convergence for the average of the iterates at a rate of
@@ A.3.2 Accelerated Generalized Gradient Method Convergence Theory
We consider accelerated methods iteratively applying the “generalized gradient step” (30) following the development of Nesterov [17] (this can also be viewed as a prox linear step [32] on the composition of the maximum function $\max\{t_{1},\dots t_{m}\}$ with the smooth mapping $y\mapsto\frac{1}{2}(h^{2}_{1}(y),\dots,h^{2}_{m}(y))$). Computing this step can be formulated as a quadratic program of dimension $m$. When $m=1$, this is exactly a (sub)gradient step on $\frac{1}{2}h^{2}$ with stepsize $\alpha$. For $m=2$, this can be computed by optimality condition $\partial\left\{\max_{i=1,2}\left\{\frac{1}{2}h^{2}_{i}(y)+h_{i}(y)g_{i}^{T}(z-y)\right\}+\frac{1}{2\alpha}\|z-y\|^{2}\right\}=0$. It has the following three cases:
  1. (a)
    If $\frac{1}{2}h^{2}_{1}(y)+h_{1}(y)g_{1}^{T}(z-y)>\frac{1}{2}h^{2}_{2}(y)+h_{2}(y)g_{2}^{T}(z-y)$, optimality condition gives $z=y-\alpha h_{1}(y)g_{1}$.
  2. (b)
    Similarly, if $\frac{1}{2}h^{2}_{1}(y)+h_{1}(y)g_{1}^{T}(z-y)<\frac{1}{2}h^{2}_{2}(y)+h_{2}(y)g_{2}^{T}(z-y)$, we obtain $z=y-\alpha h_{2}(y)g_{2}$.
  3. (c)
    If $\frac{1}{2}h^{2}_{1}(y)+h_{1}(y)g_{1}^{T}(z-y)=\frac{1}{2}h^{2}_{2}(y)+h_{2}(y)g_{2}^{T}(z-y)$, optimality condition gives $z=y-\alpha\left((1-\theta)h_{1}(y)g_{1}+\theta h_{2}(y)g_{2}\right)$, where $\theta=\frac{\frac{1}{\alpha}(h_{2}^{2}(y)-h_{1}^{2}(y))+(h_{1}(y)g_{1}-h_{2}(y)g_{2})^{T}(h_{1}(y)g_{1})}{\|h_{1}(y)g_{1}-h_{2}(y)g_{2}\|_{2}^{2}}$ is calculated by substituting $z$ into the equality condition.
Then the closed form is given by
$\mathtt{gen\mbox{-}grad}(y,\alpha)=\begin{cases}y^{(1)}&\text{ if }\frac{1}{2}h^{2}_{1}(y)+h_{1}(y)g_{1}^{T}(y^{(1)}-y)>\frac{1}{2}h^{2}_{2}(y)+h_{2}(y)g_{2}^{T}(y^{(1)}-y)\\ y^{(2)}&\text{ if }\frac{1}{2}h^{2}_{1}(y)+h_{1}(y)g_{1}^{T}(y^{(2)}-y)<\frac{1}{2}h^{2}_{2}(y)+h_{2}(y)g_{2}^{T}(y^{(2)}-y)\\ y^{(3)}&\text{ otherwise}\end{cases}$
@@ A.3.3 Level Projection Method Convergence Theory
Finally, we consider the level-projection method (29). When $m=1$, (28) with $\bar{h}=p_{*}$ is exactly a (sub)gradient step with the Polyak stepsize rule. When $m=2$, this level projection step can still be directly computed in the following three cases:
  1. (a)
    If $\frac{1}{2}h^{2}_{1}(y)+h_{1}(y)g_{1}^{T}(z-y)>\frac{1}{2}h^{2}_{2}(y)+h_{2}(y)g_{2}^{T}(z-y)$, then we need to solve this projection problem
    $\begin{cases}\min_{z}&\frac{1}{2}\|z-y\|^{2}\\ \mathrm{s.t.}&\frac{1}{2}h^{2}_{1}(y)+h_{1}(y)g_{1}^{T}(z-y)=\frac{1}{2}\bar{h}^{2}\ .\end{cases}$
    Optimality condition ensures $z-y=\lambda h_{1}(y)g_{1}$ for some $\lambda>0$. Then solving for $\lambda$ by substituting zyz-y into the constraint, we obtain $\lambda=\frac{\frac{1}{2}h^{2}_{1}(y)-\frac{1}{2}\bar{h}^{2}}{\|h_{1}(y)g_{1}\|^{2}}$.
  2. (b)
    Similarly, if $\frac{1}{2}h^{2}_{1}(y)+h_{1}(y)g_{1}^{T}(z-y)<\frac{1}{2}h^{2}_{2}(y)+h_{2}(y)g_{2}^{T}(z-y)$, the same projection calculation for $i=2$ provides $\lambda=\frac{\frac{1}{2}h^{2}_{2}(y)-\frac{1}{2}\bar{h}^{2}}{\|h_{2}(y)g_{2}\|^{2}}$.
  3. (c)
    If $\frac{1}{2}h^{2}_{1}(y)+h_{1}(y)g_{1}^{T}(z-y)=\frac{1}{2}h^{2}_{2}(y)+h_{2}(y)g_{2}^{T}(z-y)$, then we need to calculate the following projection
    $\begin{cases}\min_{z}&\frac{1}{2}\|z-y\|^{2}\\ \mathrm{s.t.}&\frac{1}{2}h^{2}_{1}(y)+h_{1}(y)g_{1}^{T}(z-y)=\frac{1}{2}\bar{h}^{2}\\ &\frac{1}{2}h^{2}_{2}(y)+h_{2}(y)g_{2}^{T}(z-y)=\frac{1}{2}\bar{h}^{2}\ .\end{cases}$
    Optimality condition ensures $z-y=\lambda_{1}h_{1}(y)g_{1}+\lambda_{2}h_{2}(y)g_{2}$ for some $\lambda_{1},\lambda_{2}>0$. Then solving for $\lambda$ by substituting zyz-y into the constraints, we obtain
The closed form is given by:
@@
Let $\mathcal{\bar{Y}}=\{y\mid\frac{1}{2}h^{2}(y)\leq\frac{1}{2}\bar{h}^{2}\}\neq\emptyset$ denote the target level set, and $\mathcal{Y}_{k}=\{y\mid m_{k}(y)\leq\frac{1}{2}\bar{h}^{2}\}\supseteq\mathcal{\bar{Y}}$ denote our model’s target level set (the containment follows from $m_{k}$ lower bounding $\frac{1}{2}h^{2}$). Each orthogonal projection step ensures $y_{k}-y_{k+1}$ is normal to $\mathcal{Y}_{k}$ at $y_{k+1}$. Consequently every $\bar{y}\in\mathcal{\bar{Y}}\subseteq\mathcal{Y}_{k}$ must have $(y_{k}-y_{k+1})^{T}(y_{k+1}-\bar{y})\geq 0$. Then
$\|y_{k+1}-\bar{y}\|^{2}$ $=\|y_{k}-\bar{y}\|^{2}-2(y_{k}-y_{k+1})^{T}(y_{k+1}-\bar{y})-\|y_{k+1}-y_{k}\|^{2}$
@@
Note that this recurrence implies $D_{k}$ is monotonically decreasing. Let $K$ denote the first index with $D_{k}\leq\bar{h}^{2}/\mu$. All $k<K$ must have a geometric decrease of $D_{k+1}\leq(1-\mu/8M^{2})D_{k}$. Thus $K\leq(8M^{2}/\mu)\log(\mu D_{0}/\bar{h}^{2})$. For all $k\geq K$, the recurrence ensures $D_{k+1}\leq D_{k}-\mu^{2}D_{k}^{2}/8M^{2}\bar{h}^{2}$, which implies $D_{k}\leq 8M^{2}\bar{h}^{2}/\mu^{2}(k-K)$.  [^8] The recurrence $D_{k+1}\leq D_{k}-\alpha D_{k}^{2}$, is equivalent to $\frac{1}{D_{k}}\leq\frac{1}{D_{k+1}}-\alpha\frac{D_{k}}{D_{k+1}}$, then by $\frac{D_{k}}{D_{k+1}}\geq 1$ and $1/D_{K}>0$, we obtain $D_{k}$. By assumption, we have $T/2>2K$ and so $T/2-K+1\geq T/4+1$. Hence recurrence ensures